OpenStudy (anonymous):

Express as a Product. ln√8

OpenStudy (anonymous):

ln sqrt(8) = 1/2 ln 8 = 1/2 ln(4 x 2) Is that what you wanted?

OpenStudy (anonymous):

yes!

OpenStudy (anonymous):

cool

OpenStudy (anonymous):

so the answer is 1/2 In 8?

OpenStudy (jdoe0001):

\(\bf ln(\sqrt{8})\implies ln\left(8^{\frac{1}{2}}\right)\implies \cfrac{1}{2}ln(8)\implies \cfrac{1}{2}ln(2^3)\implies \cfrac{3}{2}ln(2)\)